Magical Mathematics[Interesting Approach]

A tank 16m long and 23m wide contains water. How many cubic metres of water must be rushed into it to make the surface rise by 50/3 cm?

nadeemafzal111
7 Years agoGrade

1 Answer

Arun
7 Years ago
height = 16 2/3 cm = 16 + 0.66 cm ------ = 16.666 cm -> 0.166 m 
Volume of water required to raise the level = 16 * 23* 0.166 
= 16.088 m^3